OpenStudy (anonymous):

solve x+1/x+3 = x-2/x + 3/20

OpenStudy (shinalcantara):

is it \[\frac{ x-2 }{ x }+\frac{ 3 }{ 20 } or \frac{ x-2 }{ x+3 }/20?\]

OpenStudy (anonymous):

oh ya first clear the question @TheBookEater

OpenStudy (anonymous):

the first one you wrote, shinalcantara

OpenStudy (shinalcantara):

\[\frac{ x+1 }{ x+3 } = \frac{ x-2 }{ x }+\frac{ 3 }{ 20 }\]

OpenStudy (shinalcantara):

simplify the right side of the equation first by getting the least common denominator

OpenStudy (shinalcantara):